At what temperature on the Fahrenheit scale will the temperature on a centigrade scale be
three times that on a Fahrenheit scale?
1
Expert's answer
2020-06-04T09:45:18-0400

We know that the Fahrenheit scale can be converted to the centigrade scale as

"t(^\circ\mathrm{C}) = (t(^\circ\mathrm{F}) - 32)\cdot 5/9."

"t(^\circ\mathrm{C})" should be equal to "3t(^\circ\mathrm{F})" , therefore we obtain an equation

"3t(^\circ\mathrm{F}) = (t(^\circ\mathrm{F}) - 32)\cdot 5/9."

"27t(^\circ\mathrm{F}) = 5t(^\circ\mathrm{F})-160, \;\; t(^\circ\mathrm{F}) = -7\dfrac{3}{11}^\circ F \approx-7.3^\circ F."